# Beyond Reference-Based Metrics: Analyzing Behaviors of Open LLMs on Data-to-Text Generation

### v1

@@ Appendix A Quintd Data
Here we describe the data sources we include in the Quintd collection tool and the procedure of collecting the Quintd-1 benchmark. To replicate the data collection, please refer to the scripts we provide.
@@ A.1 Selection of Data Sources
We settled on the data sources described in Appendix 6. All the sources can be accessed using an API. Note that some of the APIs have access limits, either for the requests made from a single account per day, or for a number of requests from an IP address within a time window. However, these limits do not severely limit the data collection process on the scale we use here.
@@ A.2 Data Collection
Table 6 summarizes the sources of data and output types for each domain.
Domain Id Source Output type
openweather OpenWeather five-day weather forecast
gsmarena GSMArena product description
ice_hockey RapidAPI ice hockey game summary
owid OurWorldInData chart caption
wikidata Wikidata entity description
Table 6: The sources of data and output types for individual domains in Quintd.
@@ A.2.1 OpenWeather
OpenWeather (OpenWeatherMap.org) is an online service which provides global weather data via web interface and API. The API responses are in the JSON format documented at the official website. For our experiments, we used the forecast5 API which allows to download a 5-day forecast with 3-hour resolution for any location specified by its GPS coordinates.
The free tier is limited to 1,000 API calls per day, which is enough to download our whole test set in one bulk. However, at the time of experiments, the free API only allowed to download the data for the time when the request was made. At the time of writing, OpenWeather is pushing a new One Call API 3.0 which allows to download weather data for any timestamp, but only 4 days ahead (instead of 5). These restrictions somehow limit the replicability of our Quintd-1 dataset (at least with the free API), but do not limit downloading a new batch of data with a similar format.
@@ A.2.2 GSMArena
GSMArena is a website providing specifications and reviews for mobile devices. For downloading the data, we used the unofficial gsmarena-api tool which returns the data in a JSON format. Note that GSMArena imposes limitations on the number of requests per IP address, which may induce delays when downloading a larger amount of data.
To create a balanced sample, we downloaded detailed specification of 10 products from each available brand, and randomly selected 100 products for each split from the downloaded set.
@@ A.2.3 RapidAPI Ice Hockey
RapidAPI is a service which provides API access to data from multiple domains, including sport, finance, entertainment, and others. Most of the APIs are provided in a freemium mode, i.e. with a limited number of API calls per day.
For Quintd, we selected the IceHockeyAPI (popularity 9.1 / 10) which provides access to ice hockey games from world top leagues. Our choice was influenced by our own personal preferences, combined with the desire to cover a sport which has not been covered previously in sports report generation.
@@ A.2.4 OurWorldInData
OurWorldInData is a public database and web interface for data about world developments in various domains and sources. We used the official API (currently experimental) which is accessible through the Python package owid-catalog. The package allows to access individual CSV tables as Pandas dataframes.
For our data collection, we decided to limit ourselves to time series, i.e. a single column with values changing over time. Besides the simplicity of visualizing such a chart (which is used by human annotators for checking the correctness of the output), there is also a clear goal for the target chart description: describing the developments of a value over time. We also limited ourselves to the health domain. In particular, we selected the tables COVID data (columns new_cases_smoothed_per_million, new_tests_smoothed_per_thousand, people_vaccinated_per_hundred, reproduction_rate, and positive_rate) and Life expectancy data (column life_expectancy_0).
We downloaded the data for all countries with non-empty entries in the table, taking a random sample of 100 examples for each split. On model input, we formatted the data for each time series as two-column CSV, including the title, the description, and the unit for each example as a comment (#) at the beginning of the input.
@@ A.2.5 Wikidata
Wikidata is a large open-source knowledge graph, containing factual information about entities and their properties. Wikidata provides access through an official API, but we instead decided to extract our data using the wikidatasets Boschin and Bonald (2019) Python library, which provides access to preprocessed properties of entities from particular domains. It allowed us to avoid crawling and filtering the knowledge graph and its offline processing made the data collection faster. [^9] All the entities and properties are linked with an identifier to the Wikidata database, making the process also replicable through the official API.
@@
Appendix B GPT-4 Evaluation
We used the prompt in Figure 4 for instantiating the GPT-4-based metric. [^10] Note that the example in the prompt differs from the example used for human annotators (see Figure 5). We revised the example to be more instructive but we were not able to re-run the GPT-4 evaluation due to our limited budget. We ensured that the output is a valid JSON using a recently introduced parameter response_format in the OpenAI API. At the price of \$0.01 per 1k input tokens and \$0.03 per 1k generated tokens, the evaluation process cost approximately \$45 in total.
B.1 Aligning the Errors
@@
We tried to respect monotonic ordering of text spans but fell back to full-text search if the span is not found following the previous one. We consider this approach successful since matching completely failed only in a minority of cases (137 out of 6927). Based on our manual examination, these mostly include cases where GPT-4 tried to suggest a missing piece of text as an error, or did not manage to copy the input text verbatim.
Appendix C Human Evaluation
As described in §4.2, we set up the human evaluation campaign on Prolific. To make the data more accessible to the annotators, we created custom data visualizations for each domain. For the data in openweather and owid, we used interactive graphs from Highcharts.com, and we manually created the tables for other domains. You can find the full instructions for human annotators in Figure 5 and the examples of data visualizations in Appendix E.
Appendix D Metric Correlation
The Pearson correlation coefficients (§5.2) were computed using two lists (for $\mathcal{E}_{\text{hum}}$ and $\mathcal{E}_{\text{gpt}}$) as follows (note that each error category was considered separately):
  • For $r_{\text{domain}}$, we concatenated the average error counts per domain (see Table 5).
  • For $r_{\text{example}}$, we concatenated the count of errors per example.
  • For $r_{\text{token}}$, we concatenated the binary indicators marking an error per token.
@@ Appendix E Examples
Here we present an example of inputs and model outputs (along with annotations) for each domain:
  • openweather: Figure 7 (in) and Table 7 (out),
  • gsmarena: Figure 8 (in) and Table 7 (out),
  • ice_hockey: Figure 9 (in) and Table 9 (out),
  • owid: Figure 10 (in) and Table 10 (out),
  • wikidata: Figure 11 (in) and Table 11 (out).
@@ Appendix F Full Results
Here we include the tables with results for individual domains:
  • Table 5 presents the average numbers of errors per output separately for each domain (the aggregated results are in Table 5),
  • Table 5 presents the ratio of outputs containing at least one error separately for each domain (the aggregated results are in Table 5).
@@
Model
Human annotations ($\mathcal{E}_{\text{hum}}$)
GPT-4 annotations ($\mathcal{E}_{\text{gpt}}$)
Llama2
@@
The weather in Niemegk, Germany, will be cloudy and rainy for the next few days. On January 3rd, the temperature will be around 9.55°C, with a feels-like temperature of 6.59°C. The weather will be light rain, with a wind speed of 6.24 m/s and a gust of 12.1 m/s. The rain will last for 3 hours, and the pressure will be 991 hPa.
The weather in Niemegk, Germany, will be cloudy and rainy for the next few days. On January 3rd, the temperature will be around 9.55°C, with a feels-like temperature of 6.59°C. The weather will be light rain, with a wind speed of 6.24 m/s and a gust of 12.1 m/s. The rain will last for 3 hours, and the pressure will be 991 hPa.
Mistral
@@
From January 3rd to January 7th, the city of Niemegk in Germany will experience a range of weather conditions. On January 3rd, the temperature will be 9.55°C with a feels like temperature of 6.59°C. There will be light rain with a 3-hour precipitation of 1.06mm. The wind speed will be 6.24m/s with a gust of 12.1m/s. On January 4th, the temperature will be 8.2°C with a feels like temperature of 5.64°C. There will be light rain with a 3-hour precipitation of 1.94mm. The wind speed will be 4.31m/s with a gust of 11.6m/s. On January 5th, the temperature will be 6.37°C with a feels like temperature of 2.25°C. There will be light rain with a 3-hour precipitation of 1.13mm. The wind speed will be 7.07m/s with a gust of 15.51m/s. On January 6th, the temperature will be 2.17°C with a feels like temperature of -3.55°C. There will be overcast clouds with a 3-hour precipitation of 0.31mm. The wind speed will be 3.54m/s with a gust of 7.11m/s. On January 7th, the temperature will be -0.89°C with a feels like temperature of -0.89°C. There will be overcast clouds with a 3-hour precipitation of 0.31mm. The wind speed will be 0.63m/s with a gust of 0.7m/s.
From January 3rd to January 7th, the city of Niemegk in Germany will experience a range of weather conditions. On January 3rd, the temperature will be 9.55°C with a feels like temperature of 6.59°C. There will be light rain with a 3-hour precipitation of 1.06mm. The wind speed will be 6.24m/s with a gust of 12.1m/s. On January 4th, the temperature will be 8.2°C with a feels like temperature of 5.64°C. There will be light rain with a 3-hour precipitation of 1.94mm. The wind speed will be 4.31m/s with a gust of 11.6m/s. On January 5th, the temperature will be 6.37°C with a feels like temperature of 2.25°C. There will be light rain with a 3-hour precipitation of 1.13mm. The wind speed will be 7.07m/s with a gust of 15.51m/s. On January 6th, the temperature will be 2.17°C with a feels like temperature of -3.55°C. There will be overcast clouds with a 3-hour precipitation of 0.31mm. The wind speed will be 3.54m/s with a gust of 7.11m/s. On January 7th, the temperature will be -0.89°C with a feels like temperature of -0.89°C. There will be overcast clouds with a 3-hour precipitation of 0.31mm. The wind speed will be 0.63m/s with a gust of 0.7m/s.
Zephyr
@@
Over the next five days, the weather in Niemegk, Germany will be characterized by fluctuating temperatures and varying precipitation. On Thursday, January 3, light rain and a low temperature of 9.55°C can be expected, with a chance of rain continuing into Friday. Snowfall is possible on Friday, with a low temperature of -0.89°C. Saturday will be overcast with a low temperature of 0.03°C, and light rain is forecasted for Saturday evening. Sunday will be overcast as well, with a low temperature of -2.23°C. Monday will see a drop in temperatures, with a low of -4.07°C and scattered clouds. Temperatures will continue to fall throughout the week, with lows of -5.19°C and -9.51°C on Tuesday and Wednesday, respectively. Wind speeds will vary from 3.1 to 15.51 meters per second, with gusts reaching up to 15.51 meters per second on Monday.
Over the next five days, the weather in Niemegk, Germany will be characterized by fluctuating temperatures and varying precipitation. On Thursday, January 3, light rain and a low temperature of 9.55°C can be expected, with a chance of rain continuing into Friday. Snowfall is possible on Friday, with a low temperature of -0.89°C. Saturday will be overcast with a low temperature of 0.03°C, and light rain is forecasted for Saturday evening. Sunday will be overcast as well, with a low temperature of -2.23°C. Monday will see a drop in temperatures, with a low of -4.07°C and scattered clouds. Temperatures will continue to fall throughout the week, with lows of -5.19°C and -9.51°C on Tuesday and Wednesday, respectively. Wind speeds will vary from 3.1 to 15.51 meters per second, with gusts reaching up to 15.51 meters per second on Monday.
GPT-3.5
@@
Over the next five days, the weather in Niemegk, Germany is expected to be quite varied. On January 4th, there will be light rain with temperatures ranging from 2.17°C to 9.55°C. The following day, January 5th, there will be light snow with temperatures dropping to as low as -0.89°C. On January 6th, the weather will be overcast with temperatures ranging from -0.32°C to 1.69°C. The trend continues on January 7th with overcast clouds and temperatures dropping to as low as -2.51°C. Finally, on January 8th, there will be scattered clouds with temperatures ranging from -5.19°C to -3.67°C. Throughout this period, wind speeds will vary between 0.63 m/s and 8.12 m/s, and the pressure will range from 991 hPa to 1034 hPa.
Over the next five days, the weather in Niemegk, Germany is expected to be quite varied. On January 4th, there will be light rain with temperatures ranging from 2.17°C to 9.55°C. The following day, January 5th, there will be light snow with temperatures dropping to as low as -0.89°C. On January 6th, the weather will be overcast with temperatures ranging from -0.32°C to 1.69°C. The trend continues on January 7th with overcast clouds and temperatures dropping to as low as -2.51°C. Finally, on January 8th, there will be scattered clouds with temperatures ranging from -5.19°C to -3.67°C. Throughout this period, wind speeds will vary between 0.63 m/s and 8.12 m/s, and the pressure will range from 991 hPa to 1034 hPa.
Table 7: An example of the model outputs on openweather (see the input in Figure 7).
@@
Model
Human annotations ($\mathcal{E}_{\text{hum}}$)
GPT-4 annotations ($\mathcal{E}_{\text{gpt}}$)
Llama2
@@
The Gigabyte GSmart Mika M3 is a smartphone that was released in 2014. It has a 5-inch IPS LCD display with a resolution of 720x1280 pixels and a 13MP main camera. The device also has an 8MP front camera and is powered by a Mediatek MT6582 chipset with 1GB of RAM and 8GB of internal storage. It has dual SIM support and is available in Navy Blue.
The Gigabyte GSmart Mika M3 is a smartphone that was released in 2014. It has a 5-inch IPS LCD display with a resolution of 720x1280 pixels and a 13MP main camera. The device also has an 8MP front camera and is powered by a Mediatek MT6582 chipset with 1GB of RAM and 8GB of internal storage. It has dual SIM support and is available in Navy Blue.
Mistral
@@
The GSmart Mika M3 is a smartphone developed by Gigabyte. It was released in November 2014 and is powered by the Android 4.4.2 operating system. The device features a 5.0-inch IPS LCD display with a resolution of 720 x 1280 pixels, which translates to 294 pixels per inch. The smartphone is equipped with a 13-megapixel rear camera with autofocus and LED flash, as well as an 8-megapixel front-facing camera. It is powered by a Mediatek MT6582 chipset with a quad-core 1.3 GHz Cortex-A7 CPU and a Mali-400MP2 GPU. The device has 8GB of internal storage and 1GB of RAM, with the option to expand storage via a microSDHC card. The GSmart Mika M3 also features Wi-Fi, Bluetooth, GPS, and NFC connectivity, as well as a 3.5mm headphone jack and a microUSB 2.0 port for charging and data transfer. The device is available in Navy Blue and is powered by a 1900mAh Li-Ion battery, which provides up to 150 hours of stand-by time and up to 4 hours and 30 minutes of talk time.
The GSmart Mika M3 is a smartphone developed by Gigabyte. It was released in November 2014 and is powered by the Android 4.4.2 operating system. The device features a 5.0-inch IPS LCD display with a resolution of 720 x 1280 pixels, which translates to 294 pixels per inch. The smartphone is equipped with a 13-megapixel rear camera with autofocus and LED flash, as well as an 8-megapixel front-facing camera. It is powered by a Mediatek MT6582 chipset with a quad-core 1.3 GHz Cortex-A7 CPU and a Mali-400MP2 GPU. The device has 8GB of internal storage and 1GB of RAM, with the option to expand storage via a microSDHC card. The GSmart Mika M3 also features Wi-Fi, Bluetooth, GPS, and NFC connectivity, as well as a 3.5mm headphone jack and a microUSB 2.0 port for charging and data transfer. The device is available in Navy Blue and is powered by a 1900mAh Li-Ion battery, which provides up to 150 hours of stand-by time and up to 4 hours and 30 minutes of talk time.
Zephyr

### v3

@@ Appendix A Quintd Data
Here, we describe the data sources we include in the Quintd collection tool and the procedure of collecting the Quintd-1 benchmark. To replicate the data collection, please refer to the scripts we provide. [^14] https://github.com/kasnerz/quintd
@@ A.1 Selection of Data Sources
We settled on the data sources described in Appendix 6. All the sources can be accessed using an API. Note that some of the APIs have access limits, either for the requests made from a single account per day or for a number of requests from an IP address within a time window. However, these limits do not severely limit the data collection process on the scale we use here.
@@ A.2 Data Collection
Table 6 summarizes the output types for each domain.
Domain Id Output type
openweather five-day weather forecast
gsmarena product description
ice_hockey ice hockey game summary
owid chart caption
wikidata entity description
Table 6: The output types for individual domains in Quintd.
@@ A.2.1 OpenWeather
OpenWeather (OpenWeatherMap.org) is an online service that provides global weather data via web interface and API. The API responses are in the JSON format documented at the official website. For our experiments, we used the forecast5 API, which allows to download a 5-day forecast with 3-hour resolution for any location specified by its GPS coordinates.
The free tier is limited to 1,000 API calls per day, which is enough to download our whole test set in one bulk. However, at the time of experiments, the free API only allowed to download the data for the time when the request was made. At the time of writing, OpenWeather is pushing a new One Call API 3.0 which allows to download weather data for any timestamp, but only 4 days ahead (instead of 5). These restrictions somehow limit the replicability of our Quintd-1 dataset (at least with the free API) but do not limit downloading a new batch of data with a similar format.
@@ A.2.2 GSMArena
GSMArena is a website providing specifications and reviews for mobile devices. For downloading the data, we used the unofficial gsmarena-api tool, which returns the data in a JSON format. Note that GSMArena imposes limitations on the number of requests per IP address, which may induce delays when downloading a larger amount of data.
To create a balanced sample, we downloaded detailed specifications of 10 products from each available brand and randomly selected 100 products for each split from the downloaded set.
@@ A.2.3 RapidAPI Ice Hockey
RapidAPI is a service that provides API access to data from multiple domains, including sport, finance, entertainment, and others. Most APIs are provided in a freemium mode, i.e., with a limited number of daily API calls.
For Quintd, we selected the IceHockeyAPI (popularity 9.1 / 10), which provides access to ice hockey games from world top leagues. Our choice was influenced by our own personal preferences, combined with the desire to cover a sport that has not been covered previously in sports report generation.
@@ A.2.4 OurWorldInData
OurWorldInData is a public database and web interface for data about world developments in various domains and sources. We used the official API (currently experimental), which is accessible through the Python package owid-catalog. The package allows accessing individual CSV tables as Pandas dataframes.
For our data collection, we decided to limit ourselves to time series, i.e., a single column with values changing over time. Besides the simplicity of visualizing such a chart (which is used by human annotators for checking the correctness of the output), there is also a clear goal for the target chart description: describing the developments of a value over time. We also limited ourselves to the health domain. In particular, we selected the tables COVID data (columns new_cases_smoothed_per_million, new_tests_smoothed_per_thousand, people_vaccinated_per_hundred, reproduction_rate, and positive_rate) and Life expectancy data (column life_expectancy_0).
We downloaded the data for all countries with non-empty entries in the table, taking a random sample of 100 examples for each split. On model input, we formatted the data for each time series as a two-column CSV, including the title, the description, and the unit for each example as a comment (#) at the beginning of the input.
@@ A.2.5 Wikidata
Wikidata is a large open-source knowledge graph containing factual information about entities and their properties. Wikidata provides access through an official API, but we instead decided to extract our data using the wikidatasets Boschin (2019) Python library, which provides access to preprocessed properties of entities from particular domains. It allowed us to avoid crawling and filtering the knowledge graph, and its offline processing made the data collection faster. [^15] All the entities and properties are linked with an identifier to the Wikidata database, making the process also replicable through the official API.
@@
Appendix B Human Evaluation
As described in §4.1, we set up the human evaluation campaign on Prolific. To make the data more accessible to the annotators, we created custom data visualizations for each domain. For the data in openweather and owid, we used interactive graphs from Highcharts.com, and we manually created the tables for other domains. You can find the full instructions for human annotators in section 5 and the examples of data visualizations in Appendix E.
Appendix C GPT-4 Evaluation
We used the prompt in Figure 4 for instantiating the GPT-4-based metric. [^16] Note that the example in the prompt differs from the example used for human annotators (see section 5). We revised the example to be more instructive, but we were not able to re-run the GPT-4 evaluation due to our limited budget. We set the temperature to 0 to improve the replicability of the process. We ensured that the output is a valid JSON using the parameter response_format in the OpenAI API. At the price of \$0.01 per 1k input tokens and \$0.03 per 1k generated tokens, the evaluation process costs approximately \$45 in total.
C.1 Aligning the Errors
@@
We tried to respect the monotonic ordering of text spans but fell back to full-text search if the span is not found following the previous one. We consider this approach successful since matching completely failed only in a minority of cases (137 out of 6927). Based on our manual examination, these mostly include cases where GPT-4 tried to suggest a missing piece of text as an error or did not manage to copy the input text verbatim.
Appendix D Experiments with Open LLMs as Evaluators
To select the most suitable LLM for automatic evaluation, we compared correlations with human judgment of the following models:
  • GPT-4 OpenAI (2023a) used via OpenAI API (gpt-4-1106-preview),
  • GPT-3.5 OpenAI (2023b) used via OpenAI API (gpt-3.5-turbo-1106),
  • Llama-3-70B [^17] https://llama.meta.com/llama3/ running locally via Ollama in 4-bit quantization (meta-llama/Meta-Llama-3-70B).
We used all the models with the same prompts, temperature 0, and force-decoded JSON outputs. In Table 7, we show Pearson correlation coefficients computed as described in §5.2. We can see that the strongest model is GPT-4, followed by Llama-3-70B and GPT-3.5. As the gap between the models is substantial, we opted for using GPT-4 which is the strongest model.
model $r_{\text{word}}$ $r_{\text{example}}$ $r_{\text{domain}}$
GPT-4 0.26 0.52 0.93
GPT-3.5 0.07 0.33 0.82
Llama3-70B 0.09 0.44 0.92
Table 7: Pearson correlation coefficients of the model annotations as compared with human annotations (cf. §5.2).
@@ Appendix E Examples
Here, we present an example of inputs and model outputs (along with annotations) for each domain:
@@ Appendix F Full Results
Here, we include the tables with results for individual domains:
  • section 5 presents the average numbers of errors per output separately for each domain (the aggregated results are in section 5),
  • Table 14 presents the ratio of outputs containing at least one error separately for each domain (the aggregated results are in section 5).
@@
You are an expert data-to-text error annotation system. You undestand structured data and you can correcly operate with units and numerical values. You are designed to output token-level annotations in JSON.
@@
Given the data: ``` data ``` Annotate all the errors in the following text: ``` text ``` Output the errors as a JSON list "errors" in which each object contains fields "reason", "text", and "type". The value of "text" is the text of the error. The value of "reason" is the reason for the error. The value of "type" is one of 0, 1, 2, 3 based on the following list: - 0: Incorrect fact: The fact in the text contradicts the data. - 1: Not checkable: The fact in the text cannot be checked in the data. - 2: Misleading: The fact in the text is misleading in the given context. - 3: Other: The text is problematic for another reason, e.g. grammatically or stylistically incorrect, irrelevant, or repetitive. The list should be sorted by the position of the error in the text. *Example:* data: ``` [ [ "Aditi Bhagwat", "occupation", "television actor" ], [ "Aditi Bhagwat", "date of birth", "18 January 1981" ] ] ``` text: ``` Aditi Bhagwat, born on January 18, 1991, used to be a popular Indian television actor. The data comes from a knowledge graph. ``` output: ``` "errors": ["reason": "The data mentions that the actor was born on 1981", "text": "1991", "type": 0, "reason": "Misleadingly suggests that the actor is not alive", "text": "used to be", type: 2, "reason": "Popularity is not mentioned in the data", "text": "popular", type: 1, "reason", "Nationality is not mentioned in the data", "text": "Indian", type: 1, "reason": "The note is superfluous", "text": "The data comes from a knowledge graph.", type: 3] ``` Note that some details may not be mentioned in the text: do not count omissions as errors. Also do not be too strict: some facts can be less specific than in the data (rounded values, shortened or abbreviated text, etc.), do not count these as errors. If there are no errors in the text, "errors" will be an empty list.
@@
Model Human annotations ($\mathcal{E}_{\text{hum}}$) GPT-4 annotations ($\mathcal{E}_{\text{gpt}}$)
@@
Llama2 The weather in Niemegk, Germany, will be cloudy and rainy for the next few days. On January 3rd, the temperature will be around 9.55°C, with a feels-like temperature of 6.59°C. The weather will be light rain, with a wind speed of 6.24 m/s and a gust of 12.1 m/s. The rain will last for 3 hoursNC, and the pressure will be 991 hPa. The weather in Niemegk, Germany, will be cloudy and rainy for the next few days. On January 3rd, the temperature will be around 9.55°C, with a feels-like temperature of 6.59°C. The weather will be light rain, with a wind speed of 6.24 m/s and a gust of 12.1 m/s. The rain will last for 3 hoursM, and the pressure will be 991 hPa.
@@
Mistral From January 3rd to January 7th, the city of Niemegk in Germany will experience a range of weather conditions. On January 3rd, the temperature will be 9.55°C with a feels like temperature of 6.59°C. There will be light rain with a 3-hour precipitationNC of 1.06mm. The wind speed will be 6.24m/s with a gust of 12.1m/s. On January 4th, the temperature will be 8.2°C with a feels like temperature of 5.64°C. There will be light rain with a 3-hour precipitationI of 1.94mm. The wind speed will be 4.31m/s with a gust of 11.6m/s. On January 5th, the temperature will be 6.37°C with a feels like temperature of 2.25°CI. There will be light rain with a 3-hour precipitation of 1.13mmI. The wind speed will be 7.07m/s with a gust of 15.51m/sI. On January 6th, the temperature will be 2.17°CI with a feels like temperature of -3.55°CI. There will be overcast clouds with a 3-hour precipitation of 0.31mmI. The wind speed will be 3.54m/s with a gust of 7.11m/sI. On January 7th, the temperature will be - 0.89°C with a feels like temperature of -0.89°CI. There will be overcast clouds with a 3-hour precipitation of 0.31mm. The wind speed will be 0.63m/s with a gust of 0.7m/sI. From January 3rd to January 7th, the city of Niemegk in Germany will experience a range of weather conditions. On January 3rd, the temperature will be 9.55°C with a feels like temperature of 6.59°C. There will be light rain with a 3-hour precipitation of 1.06mm. The wind speed will be 6.24m/s with a gust of 12.1m/s. On January 4th, the temperature will be 8.2°C with a feels like temperature of 5.64°C. There will be light rain with a 3-hour precipitation of 1.94mm. The wind speed will be 4.31m/s with a gust of 11.6m/s. On January 5th, the temperature will be 6.37°C with a feels like temperature of 2.25°C. There will be light rain with a 3-hour precipitation of 1.13mmI. The wind speed will be 7.07m/s with a gust of 15.51m/s. On January 6th, the temperature will be 2.17°C with a feels like temperature of -3.55°C. There will be overcast clouds with a 3-hour precipitation of 0.31mmI. The wind speed will be 3.54m/sI with a gust of 7.11m/sI. On January 7th, the temperature will be -0.89°CI with a feels like temperature of -0.89°CI. There will be overcast clouds with a 3-hour precipitation of 0.31mmI. The wind speed will be 0.63m/sI with a gust of 0.7m/sI.
@@
Zephyr Over the next five days, the weather in Niemegk, Germany will be characterized by fluctuating temperatures and varying precipitation. On Thursday, January 3I, light rain and a lowI temperature of 9.55°C can be expected, with a chance of rain continuing into Friday. Snowfall is possible on Friday, with a low temperature of -0.89°C. Saturday will be overcast with a low temperature of 0.03°CI, and light rain is forecastedI for Saturday evening. Sunday will be overcast as well, with a low temperature of - 2.23°CI. Monday will see a drop in temperatures, with a low of - 4.07°CI and scattered clouds. Temperatures will continue to fall throughout the week, with lows of -5.19°C and -9.51°C on Tuesday and Wednesday, respectivelyNC. Wind speeds will vary from 3.1 to 15.51I meters per second, with gusts reaching up to 15.51I meters per second on Monday. Over the next five days, the weather in Niemegk, Germany will be characterized by fluctuating temperatures and varying precipitation. On Thursday, January 3I, light rain and a low temperature of 9.55°CI can be expected, with a chance of rain continuing into Friday. Snowfall is possible on Friday, with a low temperature of -0.89°CI. Saturday will be overcast with a low temperature of 0.03°C, and light rain is forecasted for Saturday eveningI. Sunday will be overcast as well, with a low temperature of -2.23°CI. Monday will see a drop in temperatures, with a low of -4.07°CI and scattered clouds. Temperatures will continue to fall throughout the week, with lows of -5.19°C and -9.51°C on Tuesday and WednesdayI, respectively. Wind speeds will vary from 3.1 to 15.51 meters per second, with gusts reaching up to 15.51 meters per second on MondayI.
@@
GPT-3.5 Over the next five days, the weather in Niemegk, Germany is expected to be quite varied. On January 4th, there will be light rain with temperatures ranging from 2.17°C to 9.55°CI. The following day, January 5th, there will be lightI snow with temperatures dropping to as low as -0.89°C. On January 6th, the weather will be overcast with temperatures ranging from -0.32°C to 1.69°C. The trend continues on January 7th with overcast clouds and temperatures dropping to as low as - 2.51°CI. Finally, on January 8th, there will be scattered clouds with temperatures ranging from -5.19°C to -3.67°C. Throughout this period, wind speeds will vary between 0.63 m/s and 8.12 m/s, and the pressure will range from 991I hPa to 1034 hPa. Over the next five days, the weather in Niemegk, Germany is expected to be quite varied. On January 4th, there will be light rain with temperatures ranging from 2.17°C to 9.55°CI. The following day, January 5th, there will be light snow with temperatures dropping to as low as -0.89°CI. On January 6th, the weather will be overcast with temperatures ranging from -0.32°C to 1.69°CI. The trend continues on January 7th with overcast clouds and temperatures dropping to as low as -2.51°CI. Finally, on January 8th, there will be scattered clouds with temperatures ranging from -5.19°C to -3.67°CI. Throughout this period, wind speeds will vary between 0.63 m/s and 8.12 m/sI, and the pressure will range from 991 hPaI to 1034 hPaI.
Table 8: An example of the model outputs on openweather (see the input in Table 8).
@@
Model Human annotations ($\mathcal{E}_{\text{hum}}$) GPT-4 annotations ($\mathcal{E}_{\text{gpt}}$)
@@
Llama2 The Gigabyte GSmart Mika M3 is a smartphone that was released in 2014. It has a 5-inch IPS LCD display with a resolution of 720x1280 pixels and a 13MP main camera. The device also has an 8MP front camera and is powered by a Mediatek MT6582 chipset with 1GB of RAM and 8GB of internal storage. It has dual SIM support and is available in Navy Blue. The Gigabyte GSmart Mika M3 is a smartphone that was released in 2014. It has a 5-inchI IPS LCD display with a resolution of 720x1280 pixels and a 13MP main camera. The device also has an 8MP front cameraNC and is powered by a Mediatek MT6582 chipset with 1GB of RAM and 8GB of internal storage. It has dual SIM support and is available in Navy BlueNC.
@@
Mistral The GSmart Mika M3 is a smartphone developed by Gigabyte. It was released in November 2014 and is powered by the Android 4.4.2 operating system. The device features a 5.0-inch IPS LCD display with a resolution of 720 x 1280 pixels, which translates to 294 pixels per inch. The smartphone is equipped with a 13-megapixel rear camera with autofocusNC and LED flash, as well as an 8-megapixel front-facing camera. It is powered by a Mediatek MT6582 chipset with a quad-core 1.3 GHz Cortex-A7 CPU and a Mali-400MP2 GPU. The device has 8GB of internal storage and 1GB of RAM, with the option to expand storage via a microSDHC card. The GSmart Mika M3 also features Wi-Fi, Bluetooth, GPS, and NFC connectivityI, as well as a 3.5mm headphone jack and a microUSB 2.0 port for charging and data transfer. The device is available in Navy Blue and is powered by a 1900mAh Li-Ion battery, which provides up to 150 hours of stand-by time and up to 4 hours and 30 minutes of talk time. The GSmart Mika M3 is a smartphone developed by Gigabyte. It was released in November 2014I and is powered by the Android 4.4.2 operating system. The device features a 5.0-inch IPS LCD display with a resolution of 720 x 1280 pixels, which translates to 294 pixels per inch. The smartphone is equipped with a 13-megapixel rear camera with autofocus and LED flash, as well as an 8-megapixel front-facing camera. It is powered by a Mediatek MT6582 chipset with a quad-core 1.3 GHz Cortex-A7 CPU and a Mali-400MP2 GPU. The device has 8GB of internal storage and 1GB of RAM, with the option to expand storage via a microSDHC card. The GSmart Mika M3 also features Wi-Fi, Bluetooth, GPS, and NFC connectivityI, as well as a 3.5mm headphone jack and a microUSB 2.0 port for charging and data transfer. The device is available in Navy BlueM and is powered by a 1900mAh Li-Ion battery, which provides up to 150 hours of stand-by time and up to 4 hours and 30 minutes of talk time.